# Raindrop GS: A Benchmark for 3D Gaussian Splatting under Raindrop Conditions

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) under raindrop conditions suffers from severe occlusions and optical distortions caused by raindrop contamination on the camera lens, substantially degrading reconstruction quality. Existing benchmarks typically evaluate 3DGS using synthetic raindrop images with known camera poses (constrained images), assuming ideal conditions. However, in real-world scenarios, raindrops often interfere with accurate camera pose estimation and point cloud initialization. Moreover, a significant domain gap between synthetic and real raindrops further impairs generalization. To tackle these issues, we introduce RaindropGS, a comprehensive benchmark designed to evaluate the full 3DGS pipeline—from unconstrained, raindrop-corrupted images to clear 3DGS reconstructions. Specifically, the whole benchmark pipeline consists of three parts: data preparation, data processing, and raindrop-aware 3DGS evaluation, including types of raindrop interference, camera pose estimation and point cloud initialization, single image rain removal comparison, and 3D Gaussian training comparison. First, we collect a real-world raindrop reconstruction dataset, in which each scene contains three aligned image sets: raindrop-focused, background-focused, and rain-free ground truth, enabling a comprehensive evaluation of reconstruction quality under different focus conditions. Through comprehensive experiments and analyses, we reveal critical insights into the performance limitations of existing 3DGS methods on unconstrained raindrop images and the varying impact of different pipeline components: the impact of camera focus position on 3DGS reconstruction performance, and the interference caused by inaccurate pose and point cloud initialization on reconstruction. These insights establish clear directions for developing more robust 3DGS methods under raindrop conditions.
@@ 1 Introduction
3D Gaussian Splatting (3DGS) in raindrop-affected scenes presents significant challenges, as adherent raindrops on camera lenses cause severe occlusions and optical distortions Li et al. (2024); Liu et al. (2025); Qian et al. (2024). These artifacts disrupt image correspondence, degrade the quality of camera pose estimation and point cloud initialization Zhu et al. (2023); Lin et al. (2025a), both of which are essential for achieving reliable 3DGS reconstruction under raindrop-affected inputs. Moreover, the presence of raindrops varies across views, blurring images by changing the camera focal plane You et al. (2013), introducing multi-view inconsistencies that further hinder reconstruction fidelity Petrovska & Jutzi (2025).
@@
Several recent methods Li et al. (2024); Liu et al. (2025); Qian et al. (2024) have explored 3D Gaussian Splatting under raindrop scenarios and demonstrated promising results on synthetic datasets. However, such evaluation settings are overly idealized and fail to capture the complexity and diversity of real-world conditions. To be specific, these methods typically assume the raindrop inputs are constrained images, where a clear details of both raindrops shape and background scenes, a good camera pose and point cloud initialization. However, acquiring such information from real-world raindrop-affected images is challenging Huang et al. (2025); Zhang et al. (2024). Inaccuracies in pose estimation and point cloud initialization can significantly degrade the quality of subsequent 3DGS reconstruction Wang et al. (2024); Fu et al. (2024); Huang et al. (2025). Furthermore, the substantial domain gap between synthetic and real raindrops raises concerns about generalization. Methods validated on synthetic datasets often fail to perform well when applied to real-world scenes. As illustrated in Figure 2 (a-c), these synthetic datasets are valuable but exhibit limitations, such as the same raindrop shape and position across different views.
@@
To address these issues, we introduce RaindropGS, a comprehensive benchmark for evaluating the complete raindrop 3DGS pipeline—from unconstrained, raindrop-corrupted input images to clear 3DGS reconstructions. Specifically, our pipeline consists of three stages: data preparation, data processing, and raindrop-aware 3DGS evaluation. For data preparation, we compare the effects of different types of images (raindrop-focused and background-focused) on the subsequent reconstruction process. During data processing, we evaluate the performance of camera pose estimation and point cloud initialization, as well as single-image raindrop removal algorithms. In raindrop-aware 3DGS evaluation, we consider methods that may affect the performance of real-world raindrop reconstruction, such as raindrops and point cloud Gaussian optimization.
@@
In addition, we collect a real-world 3D reconstruction dataset captured under raindrop conditions. For each scene, three aligned image sets are acquired: raindrop-focused, background-focused, and rain-free ground truth. This design enables evaluation of the full pipeline in real-world scenarios as well as under different focus conditions. As shown in Figure 2 (d), our RaindropGS dataset reflects real-world conditions, featuring multiple focus settings and a diverse range of raindrop characteristics.
@@
Using the collected dataset, we process the images (both raindrop-focused and background-focused) to obtain the corresponding rain-free images, estimated camera pose, and initialized point cloud. To analyze the impact of raindrops on the real-world dataset collection, we use COLMAP Schonberger & Frahm (2016b) and VGGT Wang et al. (2025) to estimate the camera pose and initialize the point cloud, enabling us to investigate how sequence-based and feed-forward approaches influence the performance of 3DGS methods. We include three widely used deraining methods, Uformer Wang et al. (2022), Restormer Zamir et al. (2022), and IDT Xiao et al. (2022) in the raindrop removal stage, comparing the impact of different raindrop removal methods on subsequent 3DGS reconstruction performance. For the raindrop-aware 3DGS evaluation, we integrate multiple 3DGS variants, including the original 3DGS Kerbl et al. (2023), WeatherGS Qian et al. (2024), GS-W Zhang et al. (2024), and 3DGS-MCMC Kheradmand et al. (2024), to evaluate the impact of different reconstruction strategies on raindrop-corrupted inputs. These methods are evaluated under varying pre-processing pipelines and focus conditions to assess their robustness and adaptability.
@@
  • Full 3DGS Pipeline Benchmark: We introduce the first 3DGS benchmark for raindrop-contaminated scenes, covering the complete pipeline from unconstrained, raindrop-corrupted images to the final 3D Gaussian reconstructions.
  • First Real-world Dataset: We collect a real-world 3DGS raindrop reconstruction dataset with aligned raindrop-focused, background-focused, and rain-free ground truth images, enabling comprehensive evaluation of reconstruction quality across different focus conditions.
  • Comparative Study and Insights: We validate existing 3DGS methods on our benchmark, revealing their strengths and limitations, and providing insights into future research directions.
@@ 2 Related Work
In recent years, 3DGS has emerged as a powerful technique for scene reconstruction. Unlike NeRF Mildenhall et al. (2021), it represents scenes using a sparse set of 3D Gaussians, enabling real-time rendering Bao et al. (2025). However, standard 3DGS benchmarks assume clear input views, and performance often degrades when images contain transient occlusions such as raindrops on the lens Liu et al. (2025); Qian et al. (2024); Kulhanek et al. (2024).
To address this issue, several methods  Li et al. (2024); Qian et al. (2024); Liu et al. (2025) have been developed to improve 3D reconstruction in raindrop scenes. WeatherGS Qian et al. (2024) first generates raindrop masks to identify occluded regions and then reconstructs clear scenes by excluding these areas during 3D Gaussian Splatting. Meanwhile, DerainGS Liu et al. (2025) incorporates a dedicated image enhancement module to remove raindrop artifacts and employs supervised Gaussian‑ellipsoid fitting, achieving 3D deraining in the final output. These methods are trained on synthetic raindrops and deliver strong results under the assumption of accurate camera pose estimation and reliable point cloud initialization. However, they overlook the initial disruptions that real raindrops introduce to both pose estimation and point cloud initialization, resulting in poor generalization to real-world raindrop scenarios.
@@
To mitigate lens occlusion artifacts, single-image derain methods have been extensively studied. Early works such as Raindrop Removal Network Qian et al. (2018) leverage visual attention to segment and inpaint raindrop regions, while UMAN Shao et al. (2021) extends this idea with multiscale feature fusion. More recently, transformer-based restoration models Lin et al. (2025c; 2024; b); Chen et al. (2024) (for example, Restormer Zamir et al. (2022), Uformer Wang et al. (2022), DiT Peebles & Xie (2023) and IDT Xiao et al. (2022)) demonstrate superior restoration under heavy rainfall by modeling long range dependencies. However, these methods process each image independently and do not enforce cross-view consistency, leading to reconstruction artifacts when applied as a preprocessing step for 3D reconstruction.
@@
Current 3DGS raindrop reconstruction methods focus primarily on the Gaussian fitting stage and ignore the influence of earlier steps on the training process, such as camera pose estimation and point cloud initialization. In addition, they rely on synthetic training datasets created by Blender on clear images Liu et al. (2025); Li et al. (2024), which creates a significant domain gap and prevents accurate evaluation in real-world conditions. A few real-world datasets have tried to simulate rain on camera lenses for stereo or small scale multi view setups. DerainNeRF Li et al. (2024) captures stereo pairs by spraying water onto a glass plate in front of a calibrated rig and provides binary raindrop masks. WeatherGS Qian et al. (2024) extracts key frames from publicly available rainy videos but does not supply a ground truth reference. Overall, existing datasets remain mostly synthetic and do not reflect real-world raindrop interference, and current algorithms overlook the early stages of the pipeline, making their performance evaluation under real conditions unreliable.
@@ 3 RaindropGS Benchmark and Dataset
In this section, we describe in detail the components of the 3DGS raindrop reconstruction benchmark, which consists of three parts. The first part describes the specific data collection process. The second and third parts describe the selection and evaluation of different models.
@@ 3.1 Data Preparation
Figure 2: Example of existing raindrop 3D datasets (DerainNeRF Li et al. (2024), WeatherGS  Qian et al. (2024), DerainGS Liu et al. (2025)) and our RaindropGS Dataset. As indicated by the red boxes, existing datasets exhibit the same raindrop distribution across different viewpoints; in contrast, the green boxes illustrate the diversity of raindrop distributions in our dataset. For each viewpoint, we include both raindrop-focused and background-focused images and provide corresponding clear images for 3DGS performance evaluation.
@@
Dataset Scene count
Images
GT
Camera focus
@@
To begin with, we consider a pinhole camera model focused on the background plane Hartley & Zisserman (2003). In the absence of optical distortion (e.g., caused by raindrops), all scene elements located on the focal plane would appear sharp and well-defined Hao et al. (2019). However, raindrops adhering to a thin cover glass placed directly in front of the lens act as miniature convex lenses, introducing optical distortion and causing defocus Jin et al. (2024). When background rays intersect a raindrop, they are refracted at the curved surface of the drop decided by Snell Law Born & Wolf (2013). In contrast, rays that do not encounter any raindrop travel without deviation through the imaging system to the sensor. Consequently, refracted and non-refracted rays map to spatially distinct locations on the image plane, illustrating how the presence of raindrops directly affects the imaging distortion. Furthermore, since raindrops don’t fully transmit light, the regions under raindrops exhibit localized intensity attenuation. This attenuation produces visible artifacts.
@@
During image capture (Figure 2(d)), raindrops adhering to the front glass shift the camera’s focal plane. When many raindrops lie within the depth of field, the camera focuses on them and the background becomes blurred. Conversely, if only a few raindrops fall within the focal region, the camera focuses on the background and the raindrops appear out of focus. Most synthetic datasets ignore focus variation and render both backgrounds and raindrops sharply in focus, which limits the practicality of 3DGS reconstruction on real raindrop-degraded images. The RaindropGS dataset explicitly addresses this issue by capturing each scene under both raindrop‐focused and background‐focused conditions to support more realistic 3DGS raindrop evaluation.
@@ 3.2 Data Processing
Our data processing pipeline consists of two main components: pose estimation and point cloud initialization, and single-image raindrop removal pre-processing. Unlike existing raindrop Gaussian splatting methods that assume known camera poses and accurate point clouds, our benchmark directly estimates both the camera poses and an initial point cloud from the raindrop-affected images. This approach enables us to evaluate the robustness of the subsequent 3DGS reconstruction under raindrop-affected conditions, particularly against potential errors in pose estimation and inaccuracies in the initial point cloud. To obtain a clean 3DGS reconstruction in the raindrop-aware 3DGS evaluation stage, we apply raindrop removal techniques to the multi-view raindrop images.
@@
To estimate the camera pose and initialize the point cloud from multi-view raindrop images, we employ COLMAP Schonberger & Frahm (2016b) and Visual Geometry Grounded Transformer (VGGT) Wang et al. (2025).
COLMAP is a robust tool capable of performing both Structure-from-Motion (SfM) Schonberger & Frahm (2016a) and Multi-View Stereo (MVS) Furukawa et al. (2015). We leverage SfM to estimate intrinsic and extrinsic camera parameters and MVS to generate the initial point cloud. However, raindrop interference often impedes reliable feature matching across viewpoints. This results in significant errors in estimated camera parameters and a drastic reduction in initialized point cloud density. To overcome the limitations of SfM, we employ VGGT as a comparative baseline. VGGT, a feed-forward unified method for pose estimation and point cloud generation, is more robust to raindrop interference due to its use of DINO Oquab et al. (2023).
In raindrop-focused scenes, the background is often too blurred for reliable scene initialization, causing both COLMAP Schonberger & Frahm (2016b) and VGGT Wang et al. (2025) to fail. In certain scenes, COLMAP may suffer a substantial reduction in the number of matchable camera poses due to degraded Correspondence Search Schonberger & Frahm (2016b) performance, which ultimately leads to reconstruction failure. To address cases where raindrop interference and blur reduce the initial point cloud produced by COLMAP and VGGT, we employ a random point-cloud initialization strategy. Specifically, 100,000 points are randomly initialized, matching the order of magnitude of point counts obtained from ground-truth scenes.
@@
Since traditional 3DGS methods do not incorporate raindrop removal capabilities, we employ three widely used single-image restoration models. Uformer Wang et al. (2022) applies non-overlapping window-based self-attention and a multi-scale restoration modulator, demonstrating superior capability in restoring details from raindrop-affected and blurry images. Restormer Zamir et al. (2022) leverages multi-Dconv head transposed attention and a gated-Dconv feed-forward network to restore high-quality images, while IDT Xiao et al. (2022) employs a dual Transformer with window- and spatial-based designs for rain streak and raindrop removal.
All raindrop removal methods are trained on the Raindrop Clarity dataset Jin et al. (2024) to acquire raindrop removal capabilities. Raindrop Clarity is a dataset containing both daytime and nighttime image pairs, though we only use the daytime data for training. Furthermore, Raindrop Clarity includes both background-focused and raindrop-focused image pairs, making it well-suited for our task.

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) under raindrop conditions suffers from severe occlusions and optical distortions caused by raindrop contamination on the camera lens, substantially degrading reconstruction quality. Existing benchmarks mainly rely on synthetic raindrop datasets with known camera poses (constrained images), assuming ideal conditions. However, in real-world scenarios, raindrops severely hinder accurate pose estimation and point cloud initialization, while the large domain gap between synthetic and real data further limits generalization. To tackle these issues, we introduce RaindropGS, a comprehensive benchmark designed to evaluate the full 3DGS pipeline—from unconstrained, raindrop-corrupted images to clear 3DGS reconstructions. We first collect real-world paired 3D scenes, each containing three aligned image sets: raindrop-focused, background-focused, and rain-free ground truth. These real-world data enable reliable evaluation of reconstruction quality under different focus conditions. Building on this foundation, our benchmark estimates camera poses and initializes point clouds from raindrop-corrupted images, followed by raindrop removal for 3D Gaussian optimization. This setup enables end-to-end evaluation of how errors accumulated across the 3DGS reconstruction pipeline affect the final reconstruction quality under real-world raindrop interference. Through comprehensive experiments and analyses, we reveal critical insights into the performance limitations of existing 3DGS methods on unconstrained raindrop images and quantify the impact of individual pipeline components. These insights establish clear directions for developing more robust 3DGS methods under raindrop conditions.
Figure 2: Example of existing raindrop 3D datasets (DerainNeRF [li2024derainnerf], WeatherGS  [qian2024weathergs], DerainGS [liu2025deraings]) and our RaindropGS Dataset. As indicated by the red boxes, existing datasets exhibit the same raindrop distribution across different viewpoints; in contrast, the green boxes illustrate the diversity of raindrop distributions in our dataset. For each viewpoint, we include both raindrop-focused and background-focused images and provide corresponding clear images for 3DGS performance evaluation.
@@ 1 Introduction
3D Gaussian Splatting (3DGS) in raindrop-contaminated scenes presents significant challenges [dai2025rainygs, petrovska2025impact, li2024dual], as adherent raindrops on camera lenses cause severe occlusions and optical distortions [li2024derainnerf, liu2025deraings, qian2024weathergs, chang2024uav, yu2024mip]. These artifacts disrupt image correspondence, degrade the quality of camera pose estimation and point cloud initialization [zhu2023occlusion, kotovenko2025edgs, niu2025hgsloc, wang2024real], both of which are essential for successful 3DGS reconstruction. Moreover, the presence of raindrops varies across views, blurring images by changing the camera focal plane [you2013adherent, li2024dual, li2025ntire, you2013adherent, you2015adherent], introducing multi-view inconsistencies that further hinder reconstruction fidelity [petrovska2025seeing, lyu2024rainyscape].
@@
Several recent methods [li2024derainnerf, liu2025deraings, qian2024weathergs] have explored 3D Gaussian Splatting under raindrop scenarios and demonstrated promising results on synthetic datasets. However, such evaluation settings are overly idealized and fail to capture the complexity and diversity of real-world conditions. To be specific, these methods typically assume the raindrop inputs are constrained images, where a clear details of both raindrops shape and background scenes, a good camera pose and point cloud initialization. However, acquiring such information from real-world raindrop-affected images is challenging [huang20253r, zhang2024GS-W, pan2025liberated]. Inaccuracies in pose estimation and point cloud initialization can significantly degrade the quality of subsequent 3DGS reconstruction [wang2024pfgs, fu2024colmapfree, huang20253r]. Furthermore, the substantial domain gap between synthetic and real raindrops raises concerns about generalization. Methods validated on synthetic datasets often fail to perform well when applied to real-world scenes.
@@
To address these issues, we introduce RaindropGS, a comprehensive benchmark for evaluating the complete raindrop 3DGS pipeline—from unconstrained, raindrop-corrupted input images to clear 3DGS reconstructions. As illustrated in Figure 1, our pipeline consists of three stages: data preparation, data processing, and raindrop-aware 3DGS evaluation. For data preparation, we compare the effects of different types of images (raindrop-focused and background-focused) on the subsequent reconstruction process. During data processing, we evaluate the performance of camera pose estimation and point cloud initialization, as well as single-image raindrop removal algorithms. In raindrop-aware 3DGS evaluation, we consider methods that may affect the performance of real-world raindrop reconstruction, such as raindrops and point cloud Gaussian optimization.
@@
As illustrated in Figure 2 (a-c), these synthetic datasets are valuable but exhibit limitations, such as the same raindrop shape and position across different views. To overcome these limitations, we collect a real-world 3D reconstruction dataset captured under raindrop conditions. For each scene, three aligned image sets are acquired: raindrop-focused, background-focused, and rain-free ground truth. This design enables evaluation of the full pipeline in real-world scenarios as well as under different focus conditions. As shown in Figure 2 (d), our RaindropGS dataset reflects real-world conditions, featuring multiple focus settings and a diverse range of raindrop characteristics.
@@
Using the collected dataset, we process the images (both raindrop-focused and background-focused) to obtain the corresponding rain-free images, estimated camera pose, and initialized point cloud. To analyze the impact of raindrops on the real-world dataset collection, we use COLMAP [schonberger2016colmap] and VGGT [wang2025vggt] to estimate the camera pose and initialize the point cloud, enabling us to investigate how sequence-based and feed-forward approaches influence the performance of 3DGS methods. We include three widely used deraining methods, Uformer [wang2022uformer], AtGAN [qian2018atgan], Restormer [zamir2022restormer], and IDT [xiao2022IDT] in the raindrop removal stage, comparing the impact of different raindrop removal methods on subsequent 3DGS reconstruction performance. For the raindrop-aware 3DGS evaluation, we integrate multiple 3DGS variants, including the original 3DGS [kerbl20233d], WeatherGS [qian2024weathergs], GS-W [zhang2024GS-W], and 3DGS-MCMC [kheradmand2024MCMCGS], to evaluate the impact of different reconstruction strategies on raindrop-corrupted inputs. These methods are evaluated under varying pre-processing pipelines and focus conditions to assess their robustness and adaptability.
@@
  • We introduce the first 3DGS benchmark for raindrop-contaminated scenes, covering the complete pipeline from unconstrained, raindrop-corrupted images to the final 3D Gaussian reconstructions.
  • We collect real-world paired 3D scenes, each containing three aligned image sets: raindrop-focused, background-focused, and rain-free ground truth, enabling comprehensive evaluation of reconstruction quality across different focus conditions.
  • We comprehensively validate existing 3DGS methods on our benchmark, revealing their strengths and limitations, and providing insights into future research directions.
@@ 2 Related Work
In recent years, 3DGS has emerged as a powerful technique for scene reconstruction. Unlike NeRF [mildenhall2021nerf], it represents scenes using a sparse set of 3D Gaussians, enabling real-time rendering [bao2025survey]. However, standard 3DGS benchmarks assume clear input views, and performance often degrades when images contain transient occlusions such as raindrops on the lens [liu2025deraings, qian2024weathergs, kulhanek2024wildgaussians].
To address this issue, several methods  [li2024derainnerf, qian2024weathergs, liu2025deraings] have been developed to improve 3D reconstruction in raindrop scenes. WeatherGS [qian2024weathergs] first generates raindrop masks to identify occluded regions and then reconstructs clear scenes by excluding these areas during 3D Gaussian Splatting. Meanwhile, DerainGS [liu2025deraings] incorporates a dedicated image enhancement module to remove raindrop artifacts and employs supervised Gaussian‑ellipsoid fitting, achieving 3D deraining in the final output. These methods are trained on synthetic raindrops and deliver strong results under the assumption of accurate camera pose estimation and reliable point cloud initialization. However, they overlook the initial disruptions that real raindrops introduce to both pose estimation and point cloud initialization, resulting in poor generalization to real-world raindrop scenarios.
@@
To mitigate lens occlusion artifacts, single-image derain methods have been extensively studied. Early works such as Raindrop Removal Network [qian2018attentive] leverage visual attention to segment and inpaint raindrop regions, while UMAN [shao2021uncertainty] extends this idea with multiscale feature fusion. AtGAN [qian2018atgan] removes raindrops using an attention-guided generative adversarial network that identifies raindrop regions and reconstructs the missing background. More recently, transformer-based restoration models (for example, Restormer [zamir2022restormer], Uformer [wang2022uformer] , DiT [peebles2023scalable] and IDT [xiao2022IDT]) demonstrate superior restoration under heavy rainfall by modeling long range dependencies. However, these methods process each image independently and do not enforce cross-view consistency, leading to reconstruction artifacts when applied as a preprocessing step for 3D reconstruction.
@@
Current 3DGS raindrop reconstruction methods focus primarily on the Gaussian fitting stage and ignore the influence of earlier steps on the training process, such as camera pose estimation and point cloud initialization. In addition, they rely on synthetic training datasets created by Blender on clear images [liu2025deraings, li2024derainnerf], which creates a significant domain gap and prevents accurate evaluation in real-world conditions. A few real-world datasets have tried to simulate rain on camera lenses for stereo or small scale multi view setups. DerainNeRF [li2024derainnerf] captures stereo pairs by spraying water onto a glass plate in front of a calibrated rig and provides binary raindrop masks. WeatherGS [qian2024weathergs] extracts key frames from publicly available rainy videos but does not supply a ground truth reference. Overall, existing datasets remain mostly synthetic and do not reflect real-world raindrop interference, and current algorithms overlook the early stages of the pipeline, making their performance evaluation under real conditions unreliable.
@@ 3 RaindropGS Benchmark and Dataset
Figure 1 illustrates the overall pipeline of our benchmark RaindropGS, showing the process from unconstrained raindrop‑corrupted images to clean 3D Gaussian representations. It consists of data preparation, data processing, and raindrop-aware 3DGS evaluation.
Unlike existing methods that rely on synthetic datasets for quantitative evaluation, we collect real‑world raindrop/ground‑truth image pairs, enabling realistic assessment of each stage in the 3DGS reconstruction pipeline. Our data processing employs structure-from-motion (SfM) or feed-forward methods to estimate camera poses and initialize the point cloud. The raindrop-aware 3DGS evaluation measures the robustness of existing 3DGS methods under real-world challenges, including inaccurate point cloud initialization and imperfect raindrop removal.
@@ 3.1 Data Preparation
Dataset Scene count Images (Real) GT (Real) Camera focus
@@
To begin with, we consider a pinhole camera model focused on the background plane [hartley2003pinhole]. In the absence of optical distortion (e.g., caused by raindrops), all scene elements located on the focal plane would appear sharp and well-defined [hao2019learning]. However, raindrops adhering to a thin cover glass placed directly in front of the lens act as miniature convex lenses, introducing optical distortion and causing defocus [jin2024raindrop]. When background rays intersect a raindrop, they are refracted at the curved surface of the drop decided by Snell Law [born2013principles]. In contrast, rays that do not encounter any raindrop travel without deviation through the imaging system to the sensor. Consequently, refracted and non-refracted rays map to spatially distinct locations on the image plane, illustrating how the presence of raindrops directly affects the imaging distortion. Furthermore, since raindrops don’t fully transmit light, the regions under raindrops exhibit localized intensity attenuation. This attenuation produces visible artifacts.
@@
During image capture (Figure 2(d)), raindrops adhering to the front glass shift the camera’s focal plane. When many raindrops lie within the depth of field, the camera focuses on them and the background becomes blurred. Conversely, if only a few raindrops fall within the focal region, the camera focuses on the background and the raindrops appear out of focus. Most synthetic datasets ignore focus variation and render both background and raindrops as sharply in focus, which may reduce 3DGS reconstruction accuracy on real images. The RaindropGS dataset explicitly addresses this issue by capturing each scene under both raindrop‐focused and background‐focused conditions to support more realistic 3DGS raindrop evaluation.
@@ 3.2 Data Processing
Our data processing pipeline consists of two main components: pose estimation and point cloud initialization, and single-image raindrop removal pre-processing. Unlike existing raindrop Gaussian splatting methods that assume known camera poses and accurate point clouds, our benchmark directly estimates both the camera poses and an initial point cloud from the raindrop-affected images. This approach enables us to evaluate the robustness of the subsequent 3DGS reconstruction against potential errors in pose estimation and inaccuracies in the initial point cloud. To obtain a clean 3DGS reconstruction in the raindrop-aware 3DGS evaluation stage, we apply raindrop removal techniques to the multi-view raindrop images.
3DGS 3DGS-MCMC GS-W WeatherGS
Focus Metrics Uformer Restormer IDT Uformer Restormer IDT Uformer Restormer IDT x
[…]
BG- focused PSNR $\uparrow$ 17.906 17.741 18.094 18.219 18.148 18.239 19.123 19.074 17.818 17.124
@@
To estimate the camera pose and initialize the point cloud from multi-view raindrop images, we employ COLMAP [schonberger2016colmap] and Visual Geometry Grounded Transformer (VGGT) [wang2025vggt].
COLMAP is a robust tool capable of performing both Structure-from-Motion (SfM) [schonberger2016SFM] and Multi-View Stereo (MVS) [furukawa2015MVS]. We leverage SfM to estimate intrinsic and extrinsic camera parameters and MVS to generate the initial point cloud. However, raindrop interference often impedes reliable feature matching across viewpoints. This results in significant errors in estimated camera parameters and a drastic reduction in initialized point cloud density. To overcome the limitations of SfM, we employ VGGT as a comparative baseline. VGGT, a feed-forward unified method for pose estimation and point cloud generation, is more robust to raindrop interference due to its use of DINO [oquab2023dinov2].
In raindrop-focused scenes, the background is often too blurred for reliable scene initialization, causing both COLMAP [schonberger2016colmap] and VGGT [wang2025vggt] to fail. In certain scenes, COLMAP may suffer a substantial reduction in the number of matchable camera poses due to degraded Correspondence Search [schonberger2016colmap] performance, which ultimately leads to reconstruction failure. To address cases where raindrop interference and blur reduce the initial point cloud produced by COLMAP and VGGT, we employ a random point-cloud initialization strategy. Specifically, 100,000 points are randomly initialized, matching the order of magnitude of point counts obtained from ground-truth scenes.
@@
Since traditional 3DGS methods do not incorporate raindrop removal capabilities, we employ four widely used single-image restoration models. Uformer [wang2022uformer] applies non-overlapping window-based self-attention and a multi-scale restoration modulator, demonstrating superior capability in restoring details from raindrop-affected and blurry images. Restormer [zamir2022restormer] leverages multi-Dconv head transposed attention and a gated-Dconv feed-forward network to restore high-quality images, while IDT [xiao2022IDT] employs a dual Transformer with window- and spatial-based designs for rain streak and raindrop removal. AtGAN [qian2018atgan] removes raindrops using an attention-guided generative adversarial network that identifies raindrop regions and reconstructs the missing background.
All raindrop removal methods are trained on the Raindrop Clarity dataset [jin2024raindrop] to acquire raindrop removal capabilities. Raindrop Clarity is a dataset containing both daytime and nighttime image pairs, though we only use the daytime data for training. Furthermore, Raindrop Clarity includes both background-focused and raindrop-focused image pairs, making it well-suited for our task.